# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ 5.3.2 Delve Into the Parameter and Hyperparameter Tuning Application
However, the best performances are generated when nature-based optimizations are integrated with deep ML models (see Table 9). Specifically, CSA+HKCAE [64_om2022effective], and WHO+fused CNN-BiGRU [82_kethineni2024intrusion] come up with highest accuracies (99.7%99.9%, and 99.71%, respectively) as well as f1-scores (98.2%98.9%, and 99.05, respectively). Apart from these IDSs, the human-inspired PO and CFNN-driven model [20_alghamdi2022hybrid] produces 99.46% accuracy and 99.76% f1-score. Importantly, all of these detection systems utilize distinct metaheuristics only for tuning the parameters. No other metaheuristics-assisted optimization is performed for feature selection or hyperparameter tuning.
@@
| [64_om2022effective] | UNSW-NB15 | CSA | HKCAE | 99.7 | 98.9 | [8_savanovic2023intrusion] | UNSW-NB15 | modified FOA | EL | 99.98 | 99.99 |
| BoT-IoT | 99.9 | 98.2 | [91_saheed2024voting] | UNSW-NB15 | GWO | EL | 100 | 99.745 | |||
| [82_kethineni2024intrusion] | TON_IoT | WHO | fused CNN Bi-GRU | 99.71 | 99.05 | BoT-IoT | 99.99 | 99.955 | |||
| [20_alghamdi2022hybrid] | UNSW-NB15 | PO | CFNN | 99.46 | 99.76 | [94_aljehane2024golden] | CICIDS-2017 | SSA | A-BiLSTM | 99.69 | 98.92 |
Hyperparameter Tuning Takeaway. We examine the 27 works that leverage metaheuristics algorithms for optimizing the hyperparameters of the deep learning models, including hybrid applications. The hyperparameters optimized in the papers are shown in Table LABEL:tab:hyperparameters, including the leveraged metaheuristics and ML models. From the table, it can be seen that four IDSs ([24_rahmani2024improvement, 36_aljebreen2023binary, 53_qaddoura2024evolving, 100_pingale2023remora]) utilize metaheuristics for optimizing both parameters and hyperparameters in the ML models, at a time.
Interestingly, Regarding best-performing systems, two nature-inspired and swarm-based metaheuristics are at the top. Particularly, modified FOA+EL(XGBoost) [8_savanovic2023intrusion], GWO+EL (DT,RF,KNN, and MLP) [91_saheed2024voting], and SSA+A-BiLSTM [94_aljehane2024golden] demonstrate highest accuracies (99.98%, 99.9%100%, and 99.69% respectively) and f1-scores (99.99%, 99.745%99.955%, and 98.92% respectively).
Importantly, according to Table 9, the first and second IDSs utilize the metaheuristics techniques only for hyperparameter tuning; whereas the last system simultaneously employs another nature-inspired optimizer (GJOA) to select an optimal feature set (hybrid application). Thus we observe hyperparameter tuning as a separate and hybrid application of the metaheuristics utilized in the top-performing IoT-IDSs. Another interesting finding is that the hybrid optimization technique is applied to a discriminative deep learning architecture; whereas the standalone hyperparameter tuning is employed to the ensemble classifiers.
@@
| [8_savanovic2023intrusion] | learning rate, min_child_weight, subsample, collsample_bytree, max_depth, gamma | KNN, XGBoost | modified FOA |
@@
| [17_elmasry2020evolving] | learning rate, momentum, decay, dropout rate, number of hidden layers, numbers of neurons of hidden layers, number of epochs, batch size, optimizer, initialization function, layer type, activation function | DNN, LSTM-RNN, DBN | double PSO |
@@
| [24_rahmani2024improvement] | (PT+HPT) number of hidden layers, neurons of each layer, weights | RdNN | GAO-AOA |
@@
| [25_ghasemi2024new] | Gamma (kernel coefficient parameter), C (the amount of regularization applied to the data) | SVM | GWO |
| [28_forestiero2021metaheuristic] | — | IoT2Vec | ABF |
@@
| [29_bahaa2022novel] | number of convolutional kernels, length of convolutional filter, activation functions in the convolutional layer, probability of nodes used between the convolutional and second layers, number of second-layer neurons, activation functions in the second layer, number of third-layer neurons, activation functions in the third layer, batch sample size, learning rate | CNN | APSO-WOA |
@@
| [36_aljebreen2023binary] | (PT+HPT) number of hidden nodes, input weighted, biases, C+ for minority positive instances, C- for most negative instances | CCR-ELM | SCA |
| [40_alruwaili2023red] | learning rate, number of epochs, batch size | EL | LCWOA |
| [52_pan2021lightweight] | k (Number of Neighbors), distance weight (w_k) | kNN | Compact SCA |
@@
| [53_qaddoura2024evolving] | (PT+HPT) weights, biases, regularization value, number of neurons, type of activation function | RWNN | GWO, PSO, MVO |
| [65_anushiya2023new] | batch size, learning rate | FR-CNN | GA |
| [70_bathula2022designing] | number of hidden neurons | ANN | SHO |
| [71_basheri2023quantum] | — | CRNN | HS |
@@
| [72_ramana2022wogru] | earning rate, number of hidden layers, input weights, epochs | GRU | WOA |
@@
| [74_saheed2023novel] | learning rate, input units, batch size, dropout, epochs, activation function, layers number, optimizer, units of hidden layer | DNN | modified PSO |
@@
| [78_gopalakrishnan2022new] | number of suitable hidden neurons of DNN, iterations of Adaboost, number of suitable bootstrap in the random forest | HR-OELM | AF-EFO |
@@
| [80_bajpai2024hybrid] | learning rate, sample sampling rate (subsample), maximum depth of the tree | XGBoost | WOA |
| [85_vijayan2024original] | — | EL | TuSO |
| [86_aburasain2024enhanced] | — | HDL | Enhanced BWO |
@@
| [89_elsedimy2024novel] | num_wolves, min range, max range, initial population, crossover rate, num_qubits, depth, max fun, shots | QSVM | Improved GWO (IGWO) |
| [91_saheed2024voting] | — | EL | GWO |
@@
| [94_aljehane2024golden] | learning rate, activation, epochs, dropout rate, batch size | A-BiLSTM | SSA |
| [100_pingale2023remora] | (PT+HPT) weights and training parameters | DMN | ROA |
@@
| [106_escorcia2023sea] | learning rate, dropout, batch size, epoch count, activation function | DBN | SpSO |
@@
| [111_dakic2024intrusion] | (XGBoost): learning rate, min child weight, subsample, colsample by tree, max depth, gamma (KNN): k, weights, distance | XGBoost and KNN | GSAPSO |
@@
| [112_latif2024dtl] | optimizer, activation function, dense units, dropout, fine-tune layers, epochs | EL | GA |
| [114_alamro2023modeling] | — | CNN+LSTM | ALO, FPA |
5.4 RQ4: What are open issues raised by the integration of metaheuristics with ML in IoT-IDS?
@@
Issues regarding ML methods. To improve classification performance, combining the advantages of multiple ML algorithms can be a promising approach. Though some research has already utilized this, more advanced techniques must be applied to keep the computational overhead in control, maintaining all security requirements since IoT devices can only operate with low power and limited resources. Another challenging thing is to make the proposed IDSs capable of analyzing real-time traffic. Alarmingly, most of the models in the literature are typically offline, that is, they are trained on different datasets and are tested on real-time data. Consequently, they need to retrain periodically, which is time-consuming and expensive. Considering the IoT environment, where the data are diverse and dynamically evolve over time, these traditional static IDSs are not sufficient in real-world big-data applications, especially at the enterprise level. In this case, incremental learning can be a viable solution [liang2012group], where the system dynamically learns continuously added features that were previously unknown. In [hu2013online], an online IDS is proposed for the dynamic distributed network. Specifically, at first, a local parameterized detection model is constructed in each node using the online Adaboost algorithm. Then, all of the local models are combined using PSO-based and SVM-driven algorithms to generate a global detection model, which achieves 99.99% accuracy and 0.37% FAR. Wahab et al. [wahab2022intrusion] devise a technique to adjust the size of the DNN’s hidden layers in an online manner so that the model can continuously learn and adapt new intrusions, and update predictions dynamically. The experimental evaluation states that their online DNN surpasses the static one in terms of false positives and false negatives by 6% and 4.5%, respectively.
5.5 RQ5: What are the unexplored metaheuristics optimization algorithms for IDS in IoT?
@@
Microbiology-inspired metaheuristics. These types of optimization techniques rely on the life cycle, immune system, social behavior, and collective behavior of viruses, bacteria, and other microorganisms. For example, the bacterial foraging optimization algorithm (BFOA) [guo2021survey], bacterial swarming algorithm (BSA), bacterial-GA foraging (BF), and quantum-inspired bacterial swarming optimization (QBSO) are some popular techniques of this category. Interestingly, a few metaheuristics are based on the replication and herd immunity concept of the coronavirus, known as coronavirus herd immunity optimizer (CHIO) [al2021coronavirus], and coronavirus optimization algorithm (COVIDOA). Other well-known microbiology-driven techniques are sperm swarm optimization algorithm (SSO), swine influenza models-based optimization (SIMBO), and symbiosis organisms search (SOS). No IoT-IDS is found to utilize any of these algorithms. However, a chaotic bacterial colony optimization (CBCO) technique is utilized for tuning the weights, biases, and number of neurons of the Elman recurrent neural network (ERNN), hence generalizing the model’s performance [hussan2023ddos]. As a result, the proposed IoT DDoS attack detection system surpasses other related systems. In [khayyat2023improved], an improved bacterial foraging optimization is employed for feature selection in the smart city anomaly detection system. Specifically, the classic BFOA is enhanced using the simulated annealing technique by incorporating decisions based on probability to achieve better convergence to the global optima and to handle the local extrema. However, a Bayesian optimization algorithm is applied to tune the hyperparameters of the multiplicative long short-term memory (MLSTM) model. The experimental evaluations show that IBFOA provides better classification accuracy with less computational complexity. Since these works are related to the intrusion detection system, microbiology-inspired metaheuristics can be a good choice to test in the IoT-IDS environment with diverse ML models.
@@
Chemistry-based metaheuristics. Though many physics and math-based metaheuristics are utilized for developing robust IoT-IDSs, we find no chemistry-inspired algorithms, adopting the concepts of chemical reactions and laws, such as molecular reaction, motion, radiation, etc. Kinetic gas molecules optimization (KGMO) [moein2014kgmo], artificial chemical reaction optimization algorithm (ACROA), and ions motion optimization algorithm (IMOA) are well-established optimization algorithms of this category. Intuitively, these algorithms can produce excellent performances like physics and math-inspired ones. However, these metaheuristics have been applied to other domains, such as cyber-physical systems [tavolato2020analytical], clustering and routing algorithms in WSN [asha2020efficient], clustering in big data environment [ding2020fuzzy], etc. Asha and Gowrishankar [asha2020efficient] utilize glowworm swarm optimization (GSO) and kinetic gas molecule optimization (KGMO) to increase network lifetime and number of transmissions in WSN. This hybrid energy-efficient algorithm outperforms the existing PSO-PSO-WSN and PSO-GSO-WSN. Moreover, KGMO generally offers fast convergence and is suitable for complex real-world optimization tasks. In [ding2020fuzzy], an extended CRO (real-coded CRO) is employed to find the optimal clusters for fuzzy clustering. Consequently, the false negative activities are reduced significantly compared to the earlier models (e.g., SVM, NB, DT, RF, and FCM). Besides, the accuracy and convergence speed are improved, especially in the big data environment. based on these remarkable outcomes in various domains, chemistry-based metaheuristics can be integrated into IoT-IDSs.
@@
Miscellaneous metaheuristics. Some other popular metaheuristics are inspired by the sunflower behavior to find the best orientation towards the sun (sunflower optimization (SFO) [ehteram2023application]), players’ intelligence to find the best position to score a goal (football game optimization (FGO)), optimization inspired from mother’s care for her children [matouvsova2023mother] and the trading method of the stock exchanges (exchange market algorithm (EMA) [ghorbani2014exchange]). Recently, Prashanth et al. [prasanth2024effective] have devised an efficient routing technique for wireless sensor networks through load balancing by an SFO. While compared with the existing approaches (e.g., CRCGA, GECR, OMPFM, and GADA-LEACH), the optimization demonstrates better results considering packet delivery ratio, packet loss, throughput, average residual energy, and delay. Taking the similar characteristics (like resource-constrainedness) of both WSN and IoT, this variant of optimization algorithms can be a suitable solution for IoT-IDSs.
@@
Metaheuristics and GAN-based Detection Systems. Generative Adversarial Network (GAN) is one of the most popular detection methods regarding the development of an IDS [ferdowsi2019generative]. Particularly, it facilitates synthetic data creation and better learning of the minority classes. Moreover, it can generate samples faster than DL methods, and is capable of identifying zero-day attacks in IoT since it learns from a wide range of attack scenarios [salimans2016improved]. Ferdowsi and Saad [ferdowsi2019generative] develop a GAN-based distributed IDS to identify malicious activities, independent of a centralized controlling process. They utilize different ANNs for both the generator and discriminator. Recently, Rahman et al. [rahman2024syn] propose a GAN-based NIDS, called SYN-GAN, aiming to handle the disproportion in the existing imbalance datasets and mimic the real-world network intrusion data. Notably, the model demonstrates 91%, 84%, and 100% accuracy in the UNSW-NB15, NSL-KDD, and BoT-IoT datasets, respectively. However, GAN faces difficulties while training with high dimensional data since the developed generator and discriminator are often complex and unstable [lakshminarayanan2017simple]. Incorporating metaheuristics with GAN can be a promising solution to resolve these issues since these optimizers have already proved efficient in feature selection, parameter optimization, and hyperparameter tuning phases. Researchers, focusing on implementing IoT-IDSs, can get insight from the work [anusha2024cyber, senthilkumar2024cloud], where GAN is integrated with different optimization techniques like war strategy optimization (WSO), gazelle optimization algorithm (GOA), and archerfish hunting optimizer (AHOA), etc. for parameter tuning and feature selection, resulting in an improved performance in the attack classification process. Though these works integrate GAN and metaheuristics in WSN and cloud, they do not solely focus on the IoT environment. Considering the scale, uniqueness, and difficulties of the IoT ecosystem, as well as the necessity of the metaheuristics and ML-integrated IoT-IDSs (as described in Section 1), we recommend applying GAN and these optimization techniques to develop IoT-IDSs.
6 Discussions

### v4

@@ V-C2 Delve Into the Parameter and Hyperparameter Tuning Application
However, the best performances are generated when nature-based optimizations are integrated with deep ML models (see Table IX). Specifically, CSA+HKCAE [128], and WHO+fused CNN-BiGRU [97] come up with highest accuracies (99.7%99.9%, and 99.71%, respectively) as well as f1-scores (98.2%98.9%, and 99.05, respectively). Apart from these IDSs, the human-inspired PO and CFNN-driven model [9] produces 99.46% accuracy and 99.76% f1-score. Importantly, all of these detection systems utilize distinct metaheuristics only for tuning the parameters. No other metaheuristics-assisted optimization is performed for feature selection or hyperparameter tuning.
@@
| [128] | UNSW-NB15 | CSA | HKCAE | 99.7 | 98.9 | [152] | UNSW-NB15 | modified FOA | EL | 99.98 | 99.99 |
| BoT-IoT | 99.9 | 98.2 | [145] | UNSW-NB15 | GWO | EL | 100 | 99.745 | |||
| [97] | TON_IoT | WHO | fused CNN Bi-GRU | 99.71 | 99.05 | BoT-IoT | 99.99 | 99.955 | |||
| [9] | UNSW-NB15 | PO | CFNN | 99.46 | 99.76 | [12] | CICIDS-2017 | SSA | A-BiLSTM | 99.69 | 98.92 |
Hyperparameter Tuning Takeaway. We examine the 27 works that leverage metaheuristics algorithms for optimizing the hyperparameters of the deep learning models, including hybrid applications. The hyperparameters optimized in the papers are shown in Table X, including the leveraged metaheuristics and ML models. From the table, it can be seen that four IDSs ([139, 11, 136, 133]) utilize metaheuristics for optimizing both parameters and hyperparameters in the ML models, at a time.
Interestingly, Regarding best-performing systems, two nature-inspired and swarm-based metaheuristics are at the top. Particularly, modified FOA+EL(XGBoost) [152], GWO+EL (DT,RF,KNN, and MLP) [145], and SSA+A-BiLSTM [12] demonstrate highest accuracies (99.98%, 99.9%100%, and 99.69% respectively) and f1-scores (99.99%, 99.745%99.955%, and 98.92% respectively).
Importantly, according to Table IX, the first and second IDSs utilize the metaheuristics techniques only for hyperparameter tuning; whereas the last system simultaneously employs another nature-inspired optimizer (GJOA) to select an optimal feature set (hybrid application). Thus we observe hyperparameter tuning as a separate and hybrid application of the metaheuristics utilized in the top-performing IoT-IDSs. Another interesting finding is that the hybrid optimization technique is applied to a discriminative deep learning architecture; whereas the standalone hyperparameter tuning is employed to the ensemble classifiers.
@@
| [152] | learning rate, min_child_weight, subsample, colsample_bytree, max_depth, gamma | KNN, XGBoost | modified FOA |
@@
| [54] | learning rate, momentum, decay, dropout rate, number of hidden layers, neurons, epochs, batch size, optimizer, initialization, activation | DNN, LSTM-RNN, DBN | double PSO |
@@
| [139] | (PT+HPT) number of hidden layers, neurons of each layer, weights | RdNN | GAO-AOA |
@@
| [69] | Gamma (kernel coefficient parameter), C (the amount of regularization applied to the data) | SVM | GWO |
| [62] | — | IoT2Vec | ABF |
@@
| [26] | number of convolutional kernels, length of convolutional filter, activation functions in the convolutional layer, probability of nodes used between the convolutional and second layers, number of second-layer neurons, activation functions in the second layer, number of third-layer neurons, activation functions in the third layer, batch sample size, learning rate | CNN | APSO-WOA |
@@
| [11] | (PT+HPT) number of hidden nodes, input weighted, biases, C+ for minority positive instances, C- for most negative instances | CCR-ELM | SCA |
| [16] | learning rate, number of epochs, batch size | EL | LCWOA |
| [131] | k (Number of Neighbors), distance weight (w_k) | kNN | Compact SCA |
@@
| [136] | (PT+HPT) weights, biases, regularization value, number of neurons, type of activation function | RWNN | GWO, PSO, MVO |
| [23] | batch size, learning rate | FR-CNN | GA |
| [30] | number of hidden neurons | ANN | SHO |
| [29] | — | CRNN | HS |
@@
| [140] | earning rate, number of hidden layers, input weights, epochs | GRU | WOA |
@@
| [146] | learning rate, input units, batch size, dropout, epochs, activation function, layers number, optimizer, units of hidden layer | DNN | modified PSO |
@@
| [71] | number of suitable hidden neurons of DNN, iterations of Adaboost, number of suitable bootstrap in the random forest | HR-OELM | AF-EFO |
@@
| [27] | learning rate, sample sampling rate (subsample), maximum depth of the tree | XGBoost | WOA |
| [170] | — | EL | TuSO |
| [5] | — | HDL | Enhanced BWO |
@@
| [55] | num_wolves, min range, max range, initial population, crossover rate, num_qubits, depth, max fun, shots | QSVM | Improved GWO (IGWO) |
| [145] | — | EL | GWO |
@@
| [12] | learning rate, activation, epochs, dropout rate, batch size | A-BiLSTM | SSA |
| [133] | (PT+HPT) weights and training parameters | DMN | ROA |
@@
| [56] | learning rate, dropout, batch size, epoch count, activation function | DBN | SpSO |
@@
| [40] | (XGBoost): learning rate, min child weight, subsample, colsample by tree, max depth, gamma (KNN): k, weights, distance | XGBoost and KNN | GSAPSO |
@@
| [106] | optimizer, activation function, dense units, dropout, fine-tune layers, epochs | EL | GA |
| [8] | — | CNN+LSTM | ALO, FPA |
V-D RQ4: What are open issues raised by the integration of metaheuristics with ML in IoT-IDS?
@@
Issues regarding ML methods. To improve classification performance, combining the advantages of multiple ML algorithms can be a promising approach. Though some research has already utilized this, more advanced techniques must be applied to keep the computational overhead in control, maintaining all security requirements since IoT devices can only operate with low power and limited resources. Another challenging thing is to make the proposed IDSs capable of analyzing real-time traffic. Alarmingly, most of the models in the literature are typically offline, that is, they are trained on different datasets and are tested on real-time data. Consequently, they need to retrain periodically, which is time-consuming and expensive. Considering the IoT environment, where the data are diverse and dynamically evolve over time, these traditional static IDSs are not sufficient in real-world big-data applications, especially at the enterprise level. In this case, incremental learning can be a viable solution [111], where the system dynamically learns continuously added features that were previously unknown. In [81], an online IDS is proposed for the dynamic distributed network. Specifically, at first, a local parameterized detection model is constructed in each node using the online Adaboost algorithm. Then, all of the local models are combined using PSO-based and SVM-driven algorithms to generate a global detection model, which achieves 99.99% accuracy and 0.37% FAR. Wahab et al. [172] devise a technique to adjust the size of the DNN’s hidden layers in an online manner so that the model can continuously learn and adapt new intrusions, and update predictions dynamically. The experimental evaluation states that their online DNN surpasses the static one in terms of false positives and false negatives by 6% and 4.5%, respectively.
V-E RQ5: What are the unexplored metaheuristics optimization algorithms for IDS in IoT?
@@
Microbiology-inspired metaheuristics. These types of optimization techniques rely on the life cycle, immune system, social behavior, and collective behavior of viruses, bacteria, and other microorganisms. For example, the bacterial foraging optimization algorithm (BFOA) [72], bacterial swarming algorithm (BSA), bacterial-GA foraging (BF), and quantum-inspired bacterial swarming optimization (QBSO) are some popular techniques of this category. Interestingly, a few metaheuristics are based on the replication and herd immunity concept of the coronavirus, known as coronavirus herd immunity optimizer (CHIO) [6], and coronavirus optimization algorithm (COVIDOA). Other well-known microbiology-driven techniques are sperm swarm optimization algorithm (SSO), swine influenza models-based optimization (SIMBO), and symbiosis organisms search (SOS). No IoT-IDS is found to utilize any of these algorithms. However, a chaotic bacterial colony optimization (CBCO) technique is utilized for tuning the weights, biases, and number of neurons of the Elman recurrent neural network (ERNN), hence generalizing the model’s performance [84]. As a result, the proposed IoT DDoS attack detection system surpasses other related systems. In [99], an improved bacterial foraging optimization is employed for feature selection in the smart city anomaly detection system. Specifically, the classic BFOA is enhanced using the simulated annealing technique by incorporating decisions based on probability to achieve better convergence to the global optima and to handle the local extrema. However, a Bayesian optimization algorithm is applied to tune the hyperparameters of the multiplicative long short-term memory (MLSTM) model. The experimental evaluations show that IBFOA provides better classification accuracy with less computational complexity. Since these works are related to the intrusion detection system, microbiology-inspired metaheuristics can be a good choice to test in the IoT-IDS environment with diverse ML models.
@@
Chemistry-based metaheuristics. Though many physics and math-based metaheuristics are utilized for developing robust IoT-IDSs, we find no chemistry-inspired algorithms, adopting the concepts of chemical reactions and laws, such as molecular reaction, motion, radiation, etc. Kinetic gas molecules optimization (KGMO) [122], artificial chemical reaction optimization algorithm (ACROA), and ions motion optimization algorithm (IMOA) are well-established optimization algorithms of this category. Intuitively, these algorithms can produce excellent performances like physics and math-inspired ones. However, these metaheuristics have been applied to other domains, such as cyber-physical systems [163], clustering and routing algorithms in WSN [25], clustering in big data environment [47], etc. Asha and Gowrishankar [25] utilize glowworm swarm optimization (GSO) and kinetic gas molecule optimization (KGMO) to increase network lifetime and number of transmissions in WSN. This hybrid energy-efficient algorithm outperforms the existing PSO-PSO-WSN and PSO-GSO-WSN. Moreover, KGMO generally offers fast convergence and is suitable for complex real-world optimization tasks. In [47], an extended CRO (real-coded CRO) is employed to find the optimal clusters for fuzzy clustering. Consequently, the false negative activities are reduced significantly compared to the earlier models (e.g., SVM, NB, DT, RF, and FCM). Besides, the accuracy and convergence speed are improved, especially in the big data environment. based on these remarkable outcomes in various domains, chemistry-based metaheuristics can be integrated into IoT-IDSs.
@@
Miscellaneous metaheuristics. Some other popular metaheuristics are inspired by the sunflower behavior to find the best orientation towards the sun (sunflower optimization (SFO) [51]), players’ intelligence to find the best position to score a goal (football game optimization (FGO)), optimization inspired from mother’s care for her children [117] and the trading method of the stock exchanges (exchange market algorithm (EMA) [70]). Recently, Prashanth et al. [135] have devised an efficient routing technique for wireless sensor networks through load balancing by an SFO. While compared with the existing approaches (e.g., CRCGA, GECR, OMPFM, and GADA-LEACH), the optimization demonstrates better results considering packet delivery ratio, packet loss, throughput, average residual energy, and delay. Taking the similar characteristics (like resource-constrainedness) of both WSN and IoT, this variant of optimization algorithms can be a suitable solution for IoT-IDSs.
@@
Metaheuristics and GAN-based Detection Systems. Generative Adversarial Network (GAN) is one of the most popular detection methods regarding the development of an IDS [61]. Particularly, it facilitates synthetic data creation and better learning of the minority classes. Moreover, it can generate samples faster than DL methods, and is capable of identifying zero-day attacks in IoT since it learns from a wide range of attack scenarios [148]. Ferdowsi and Saad [61] develop a GAN-based distributed IDS to identify malicious activities, independent of a centralized controlling process. They utilize different ANNs for both the generator and discriminator. Recently, Rahman et al. [138] propose a GAN-based NIDS, called SYN-GAN, aiming to handle the disproportion in the existing imbalance datasets and mimic the real-world network intrusion data. Notably, the model demonstrates 91%, 84%, and 100% accuracy in the UNSW-NB15, NSL-KDD, and BoT-IoT datasets, respectively. However, GAN faces difficulties while training with high dimensional data since the developed generator and discriminator are often complex and unstable [105]. Incorporating metaheuristics with GAN can be a promising solution to resolve these issues since these optimizers have already proved efficient in feature selection, parameter optimization, and hyperparameter tuning phases. Researchers, focusing on implementing IoT-IDSs, can get insight from the work [22, 153], where GAN is integrated with different optimization techniques like war strategy optimization (WSO), gazelle optimization algorithm (GOA), and archerfish hunting optimizer (AHOA), etc. for parameter tuning and feature selection, resulting in an improved performance in the attack classification process. Though these works integrate GAN and metaheuristics in WSN and cloud, they do not solely focus on the IoT environment. Considering the scale, uniqueness, and difficulties of the IoT ecosystem, as well as the necessity of the metaheuristics and ML-integrated IoT-IDSs (as described in Section I), we recommend applying GAN and these optimization techniques to develop IoT-IDSs.
VI Discussions